OpenStudy (anonymous):

Solve for x. 7(3 - x) = 8(4 - 2x)

OpenStudy (anonymous):

7(3-x)=8(4-2x) 21-7x=32-16x 21-7x+16x=32 21+9x=32 9x=32-21 9x=11 x=11/9